Exercise 3.2: Single Unknown 2. Let the demand and supply function be as follows: a) Q_d= 51 – 3P b) Q_d=30-2P Q_s=6P-10 Q_s=-6+5P Find P* and Q* by elimination of variables. (Use fractions rather than decimals) 4. If (b + d) =0 in the linear market model, can an equilibrium solution be found by using (3.4) and (3.5)? Why or why not?
Answer #1

As we equate respective demand functions and supply functions we get from a) Q_d= 51 – 3P and Q_s= 6P -10 as Q_d=Q_s therefore optimal P and Q will be respectively P=61/9 and Q=92/3.

B)here Q_d=30-2P andQ_s=-6+5P so the P=36/7 andQ=138/7.

Part 4. Has insufficient information available in the question.